OpenStudy (anonymous):

Two cars leave Denver at the same time and travel in opposite directions. One car travels 10 mi/h faster than the other car. The cars are 500 mi apart in 5 h. How fast is each car traveling?

OpenStudy (anonymous):

ion

OpenStudy (anonymous):

time = 500 /2x+10 =5 10x+50=500 10x= 450 x=45 speed of one car The other car travels at 45+10 = 55 mph